OpenStudy (anonymous):

factor completely. 64y^2 - 80y + 25

OpenStudy (anonymous):

This one is apefect square trinomial since 2*8*5=80, so it factors into \[(8y-5)^2\]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

This comes from\[(a-b)^2=a^2-2ab+b^2\]In your problem a=8y and b=5

OpenStudy (anonymous):

thank you ur really good at this

OpenStudy (anonymous):

ur welcome :})